Define flux . Calculate the flux if 5 mole of glucose is passed through an area of 0.1Cm² in 2.0 seconds .
1
Expert's answer
2019-11-01T11:24:05-0400

"J = \frac{n}{S*t} =\frac{5 mol}{0.1cm^2 *2s} = 25\frac{mol}{cm^2 *s}"